OpenStudy (fanduekisses):

What is the derivative of f(x)= ln(e^-x^2)

OpenStudy (fanduekisses):

\[f(x)=\ln(e^{-x^{2}})\]

OpenStudy (anonymous):

You have to use the chain rule. first the derivative of ln(g(x)) = (1/g(x))*g'(x) right?

OpenStudy (anonymous):

say g(x) = exp(-x^2)

OpenStudy (anonymous):

now what is the derivative of e^h(x)? it is e^h(x) * h'(x)

OpenStudy (anonymous):

say h(x) = -x^2

OpenStudy (anonymous):

what is the derivative of h(x)=-x^2? it will be -2x

OpenStudy (anonymous):

now plug it all in

OpenStudy (anonymous):

derivative of ln(e^(-x^2)) = (1/(e^(-x^2)))*(e^(-x^2))*(-2x)

OpenStudy (fanduekisses):

oh ok so first the derivative of ln(x)= 1/x so that's why 1/e^-x^2 right ...

OpenStudy (fanduekisses):

and then I know that the derivative of e^x is simply e^x unless it's e^(fx) then it is e^(fx) * f'(x)

OpenStudy (fanduekisses):

I'm getting it now lol ^_^ thanks

OpenStudy (anonymous):

that's right!

OpenStudy (anonymous):

you have to split all these functions and recognize their derivatives

OpenStudy (anonymous):

then you just "chain" it all together

OpenStudy (anonymous):

;)